Let y vary directly as the square root of x and inversely as z2 . When y is tripled and z
quadrupled, x is multiplied by k . Determine the value of k .

Grade:11

1 Answers

Vikas TU
7 years ago
If  y vary directly as the square root of x and inversely as z2 then,
y = c*root(x)
and
y = c/z^2
and also 
root(x) = 1/z^2...........(1)
When y is tripled,
3y = c*root(kx)
and
3y = c/(4z)^2
root(kx) = 1/16z^2............(2)
solving eqns. (1) and (2) we get,
k = 1/256